A financial planner has just completed an analysis of Anabell Snyder’s fixed-income holdings. The financial planner has determined each of Anabell’s after-tax yields, but is cautioning Anabell that the tax implications of her holdings might change if Congress alters marginal tax rates. Based on the following after-tax yields, which of these bonds would offer the greatest after-tax return if Anabell’s federal marginal tax bracket increased from 24 percent to 32 percent, while her state marginal bracket remained at 4.5 percent?

•A corporate bond with a 5.1 percent after-tax return.

•An out-of-state municipal bond with a 5.0 percent after-tax return.

•An in-state municipal bond with a 4.8 percent after-tax return.
Answer #1

Please do Upvote if you are served. Feel free to reach out in the comments

Cheers!!!

Answer:

Given the increase in marginal tax braxket change, Lets calculate the yield of each of the given bonds

=> Corporate bond Tax equivalent yields = 5.1% / (1 - 30%) = 7.285% ( Fed marginal tax bracket for corporate andout-of state municipal bonds)

=> Out-of-state municipal bond Tax equivalent yields = 5.0% / ( 1 -30%) = 6.857%

=> In-state municipal bond Tax equivalent Yields = 4.8% / (1 - 4.5%) = 5.026%

Therefore Corporate bonds offer the greatest afer-tax returns